OpenStudy (shes.radical):

A student is assessing the correlation between the number of workers in a factory and the number of units of output produced daily. The table below shows the data: Number of workers (x) 0 10 20 30 40 50 60 70 80 90 Number of units of output (y) 2 52 102 152 202 252 302 352 402 452 Part A: Is there any correlation between the number of workers in a factory and the number of units of output produced daily? Justify your answer. Part B: Write a function which best fits the data. Part C: What does the slope and y-intercept of the plot indicate?

OpenStudy (shes.radical):

My Answer For Part A: Yes, The more number of workers, the more units of output. That's all I got. And I think that's wrong.

OpenStudy (shes.radical):

@peachpi ..

OpenStudy (anonymous):

yes there is a correlation. If you graph it on a calculator, you will see there is a positive correlation. With the numbers in your calculator, you can run the LinReg to find the line of best fit.

OpenStudy (shes.radical):

LinReg?

OpenStudy (anonymous):

do you have a graphing calculator?

OpenStudy (shes.radical):

i do not..

OpenStudy (anonymous):

did that come up with the numbers I put in?

OpenStudy (anonymous):

Alright I'm making this too complicated. There's a way to use the calculator to find the equation of a line that you plotted. But with the numbers this gives you, it is linear, which means you can find the slope yourself by doing (y2-y1)/(x2-x1)

OpenStudy (shes.radical):

i just used the calculator..wow that thing is so helpful, and so are you, thank youuu sooooo very much

OpenStudy (anonymous):

you will learn how to use it more efficiently if you take statistics :). It's great to use

OpenStudy (shes.radical):

but wait..what about Part C?

OpenStudy (anonymous):

the y-intercept is when x is at 0. if you used the calculator I sent you, you should have gotten the equation y=5x+2. also you can see from the data that when x is 0, y is 2. So 2 is the y-intercept, and the slope is 5. That tells you that when there are 0 workers, there are 2 units of output

OpenStudy (anonymous):

you put in y=5x+2 and not y=2x+5 right?. the site gives it backwards

OpenStudy (shes.radical):

yeah i accidently put in y=2x+5 instead of y=2+5x xD i got it though, thats right. but Part C..i still dont get that part .-.

OpenStudy (shes.radical):

i think: y= the number of units slope= .....idk..

OpenStudy (anonymous):

slope is: number of units/ number of workers so if the slope is 5 then wouldn't that mean there are 5 units output for every worker?

OpenStudy (shes.radical):

but the slope is 2+5

OpenStudy (shes.radical):

er, thats what the website gave me.....y=2+5x

OpenStudy (anonymous):

y=2+5x is the same as y=5x+2 y=mx+b is the slope intercept form where m=slope, so the slope would be 5

OpenStudy (shes.radical):

ohhhhhhhhhh, so then slope represents the number of unit outputs per worker..right?

OpenStudy (anonymous):

yep!

OpenStudy (shes.radical):

too bad i can only give you 1 medal Dx you're a life saver!! thank you so much

OpenStudy (anonymous):

haha np! glad I could help

OpenStudy (shes.radical):

i have 1 more question..er half of it since i answered the first part..up for it? @helpppppppppp

OpenStudy (anonymous):

i'll try my best :)

OpenStudy (shes.radical):

cool ima post it separately though, ill tag you

OpenStudy (anonymous):

ok